OpenStudy (anonymous):

derive the value of acceleration due to gravity i know it is of physics but i need it urgently please help me any1

OpenStudy (vishweshshrimali5):

use g = GMm/r^2

OpenStudy (turingtest):

derive it from what, the general formula \[F_g=Gm_1m_2/r^2\] ? you mean F= vishweshshrimali...

OpenStudy (vishweshshrimali5):

yes !! sorry the correct formula is g = GM/r^2

OpenStudy (turingtest):

think of a falling object \[mg=Gm_em/r^2\] the m's of the falling object cancel giving \[g=Gm_e/r_e^2\] Plug in the values for the constant G=6.67x10^(-11), mass of the earth, and radius of the earth and you should get 9.8m/s^2

OpenStudy (anonymous):

wahe guru what is confusion now ?

OpenStudy (turingtest):

g=(6.67e-11)(5.9742e24)/(6.378e6)^2=9.795... about 9.8m/s^2